I realy need help with this problem:
4x-2x+4=x+10
don't understand this at all

Answer by stanbon(75887)

4x-2x+4=x+10
-------
Combine the "like" terms on the left side:
2x+4 = x+10
Subtract x from both sides:
x+4 = 10
Subtract 4 from both sides:
x = 6
===========
Cheers,
Stan H.